OpenStudy (anonymous):

solve for x: sec(x) * cos ((pi/2) - x) = -1

OpenStudy (anonymous):

Domain is [0,2pi)

OpenStudy (anonymous):

There exists a trig identity such that you can say: cos((pi/2)-x) = cos(pi/2)cos(x) + sin(pi/2)sin(x) cos(pi/2) = 0 sin(pi/2) = 1 So now our equation ends up being sec(x)*(0*cos(x) + sin(x)) = -1 that simplifies to: sec(x)*sin(x) = -1 That should be enough for you to figure out the rest. Lemme know if you need anymore help.

OpenStudy (mertsj):

\[\cos (\frac{\pi}{2}-x)=\sin x\]